1. {3(7 - 9) + 16}^3 ÷ (-5) - 2^4: 1. {3(7 - 9) + 16}^3 ÷ (-5) - 2^4
Answer by Alan3354(1926)

{3(7 - 9) + 16}^3 ÷ (-5) - 2^4
= (3(-2 + 16)^3/((-5) - 16) \\\ It's not clear what's meant to be in the denominator. Use more parentheses.
= (3*14)^3/(-21)
= 42^3/(-21)
= 42*42^2/(-21) A shortcut
= -2*42^2
= -3528